# Measuring Casimir Force Across a Superconducting Transition

### v1

@@ Acknowledgements.
where $\eta=\hbar\gamma/2\Delta(T)$, $t_{\rm red}=k_{B}T/\hbar\gamma$, and $E(x)=\sqrt{x^{2}+1/(4\eta^{2})}$. It is noted that $10<\tilde{\omega}_{s}(T)<$ 1, implying $\bar{\Omega}(T)<\Omega$. Equations S.44 and S.58 show that $\epsilon_{\rm BCS}$ exhibits a $1/\xi^{2}$ singularity as $\xi\rightarrow 0$, similar to the plasma model Equation S.57, which results in the following expression of the TE reflection coefficient of a BCS superconductor:
[…]
@@
Plasma-BCS approach: in this approach, for $T>T_{c}$ we use the plasma-model formula for $r_{\rm TE}(0,k_{\bot})$ (Equation (S.56)), and for $T<T_{c}$ the BCS expression Equation (S.61) is used. The plasma frequency $\Omega=5.33$ $\hbar$ used corresponds to the optical response of NbTiN in the infrared region. As depicted in panel (c) of Fig. (S23) this approach does predict a jump in $\Delta P^{\prime}$, consistent with our experimental observations.
@@
Fig. S24(a) and (b) present the experimental data alongside theoretical predictions from the Drude-BCS and Plasma-BCS approaches for the variations in pressure gradient ($\Delta P^{\prime}$) and membrane frequency shift $\Delta\omega^{2}$, respectively. The Drude-BCS approach is clearly inconsistent with the data, whereas the Plasma-BCS approach predicts jumps in $\Delta P^{\prime}$ and $\Delta\omega^{2}$ that are approximately half the experimental values, suggesting its potential to explain the observed phenomena. It is noted that the Plasma-BCS approach can fully reproduce the experimental data with a plasma frequency of $\Omega=12.7$ $\hbar$. As our current understanding of the plasma approach to the Casimir effect provides no strong justification for limiting $\Omega$ to the material’s infrared optical response, the potential for employing a different value cannot be dismissed.
Our observation of a sudden resonance frequency shift during the membrane’s superconducting transition, if attributable to the Casimir force, would be the first experimental evidence of the superconducting transition’s influence on the Casimir effect. However, interpreting the data in this manner reveals two unexpected features. First, the sign of the frequency shift indicates a decrease in the Casimir pressure gradient during the transition, which contradicts the intuition that a superconducting surface, acting as a near-ideal mirror, should increase the Casimir force. Second, and perhaps more significantly, the abrupt frequency shift at the critical temperature seems to be inconsistent with the second-order nature of the superconducting transition. Since the Casimir force is the negative derivative of the Helmholtz free energy ${\cal F}$ of the system with respect to separation [38], $F_{\rm Cas}=-\partial{\cal F}/\partial d$, $F_{\rm cas}$ is expected to be continuous across the transition. Furthermore, the smooth confluence of the BCS permittivity with the Drude permittivity at the superconducting transition temperature provides another reason to expect a continuous change in the Casimir force. The Drude-BCS theory of the Casimir effect in superconductors, as outlined in [14], cannot replicate the observed jump in the Casimir pressure, predicting instead a smooth increase in Casimir pressure. Conversely, the Plasma-BCS approach semi-quantitatively reproduces the observed behavior. The established agreement of the plasma approach with room-temperature Casimir experiments [39, 40, 79, 81, 90] suggests its expected applicability to the normal state of the membrane.
Figure S24: Comparison of the experimental data and the theoretical analysis, on the abrupt changes in the (a) pressure gradient, and in the (b) membrane frequency square shift. The plasma-BCS approach seems capable of explaining the abrupt changes occurring at $T_{C}$.
@@
This work $709\times 709$ $4.9\times 10^{-7}$ 190 $\bf 4.89117\times 10^{-7}$ 2025
@@
Figure S28: Schematic of the Casimir cavity realized in this paper is presented at realistic scales. The scale drawing, representing the fabricated devices, exhibits extraordinary aspect ratios: a ratio of 4574 between its lateral dimensions and the resonator thickness, and a ratio of 3732 between its lateral dimensions and the vacuum gap.
Table S4: Casimir experiments with sphere-plate configuration. The effective area of the sphere-plate configuration is defined as $A_{\rm eff}=R/d_{\rm min}$, where $R$ is the radius of the sphere, and $d_{\rm min}$ is the minimum separation. The ideal Casimir force calculated in this table is assuming the configurations are made of perfectly conducting plates, for a better comparison of their geometrical sizes, regardless of the materials.

### v2

@@ Acknowledgements.
where $\eta=\hbar\gamma/2\Delta(T)$, $t_{\rm red}=k_{B}T/\hbar\gamma$, and $E(x)=\sqrt{x^{2}+1/(4\eta^{2})}$. It is noted that $0<\tilde{\omega}_{s}(T)<1$, implying $\bar{\Omega}(T)<\Omega$. Equations S.56 and S.70 show that $\epsilon_{\rm BCS}$ exhibits a $1/\xi^{2}$ singularity as $\xi\rightarrow 0$, similar to the plasma model Equation S.69, which results in the following expression of the TE reflection coefficient of a BCS superconductor:
$r_{\rm TE}(0,k_{\bot})=\frac{k_{\bot}-\sqrt{\bar{\Omega}(T)^{2}/c^{2}+k_{\bot}^{2}}}{k_{\bot}+\sqrt{\bar{\Omega}(T)^{2}/c^{2}+k_{\bot}^{2}}},\quad\quad\text{(BCS model for $T<T_{c}$)}~.$ (S.73)
@@
[…]
@@
[…]
Figure S26: Comparison of the experimental data and the theoretical analysis, on the abrupt changes in the panel (a) pressure gradient, in panel (b) membrane angular frequency square shift, and in panel (c) membrane linear frequency shift. The plasma-BCS approach seems capable of explaining the abrupt changes occurring at $T_{C}$.
@@
In this section, we showed the mechanical quality factors of small- and big-gap membranes as a function of temperature, in Figure S27 (over a large temperature range) and Figure S28 (across $T_{C}$).
@@
This work $709\times 709$ $4.9\times 10^{-7}$ 190 $4.89117\times 10^{-7}$ 2025
@@
Table S4: Casimir experiments with sphere-plate configuration. The effective area of the sphere-plate configuration is defined as $A_{\rm eff}=R/d_{\rm min}$, where $R$ is the radius of the sphere, and $d_{\rm min}$ is the minimum separation. The Casimir force calculated in this table is assuming the configurations are made of perfectly conducting plates, for a better comparison of their geometrical sizes, regardless of the materials.